find the limit lim x→0 x^2 cos 1/x.


1
Expert's answer
2022-01-13T16:55:55-0500

Solution:

Given "\varepsilon>0" , choose "\delta=\sqrt{\varepsilon}"

Now, if "x" is chosen so that "0<|x-0|<\delta" ,

then

"\begin{aligned}

\left|x^{2} \cos \left(\frac{1}{x}\right)-0\right| &=\left|x^{2}\right|\left|\cos \left(\frac{1}{x}\right)\right| \\

&=x^{2}\left|\cos \left(\frac{1}{x}\right)\right| \\

& \leq x^{2}(1) \\

&<(\sqrt{\varepsilon})^{2}=\varepsilon

\end{aligned}"

That is: if "0<|x-0|<\delta" , then "\left|x^{2} \cos \left(\frac{1}{x}\right)-0\right|<\varepsilon"

So, by definition of limit, "\lim _{x \rightarrow 0}\left(x^{2} \cos \left(\frac{1}{x}\right)\right)=0."